OpenStudy (anonymous):

Determine the ratio and write it in lowest terms. 4 hours to 90 minutes

OpenStudy (anonymous):

first covert hours to minutes, so the units are the same

OpenStudy (anonymous):

one hour is 60 minutes, so 4 hours is \(4\times 60=2400\) minutes then compute \[\frac{2400}{90}\]

OpenStudy (mathstudent55):

You need to use the smae units. Since you have hours and minutes, convert hours to minutes first. Then you can write the ratio as (minutes in 4 hours)/(minutes in 90 minutes) Then rduce the fraction.

OpenStudy (anonymous):

finally reduce the fraction to lowest terms. you can start with \(\frac{2400}{90}=\frac{240}{9}\) and reduce further

OpenStudy (anonymous):

45/2

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

sorry no

OpenStudy (anonymous):

8/3

OpenStudy (anonymous):

divide top and bottom by 3

OpenStudy (anonymous):

whoa i made mistake sorry

OpenStudy (anonymous):

you are right,, ignore me, it is \[\frac{8}{3}\]

OpenStudy (anonymous):

lol k

OpenStudy (anonymous):

sorry about that, didn't mean to confuse you

OpenStudy (anonymous):

tha'ts alright

OpenStudy (mathstudent55):

240/90 = 8/3